# Excluding cosmetic surgeries on hyperbolic 3-manifolds

### v2

@@ 2. $\>\>(2)$
$\tfrac{1}{2}\operatorname{span}(\Delta_{K})\leq g(K).$
If $g_{T}(D)<\operatorname{span}\Delta_{K}(\tfrac{1}{2}\operatorname{span}\Delta_{K}-2)$, then Lowrance’s theorem [35] implies $th(K)<2g(K)(g(K)-2)$, hence Hanselman’s Theorem 2.1 implies $K$ cannot have purely cosmetic surgeries.
If $g_{T}(D)=0$, then $K$ is alternating [9, Corollary 4.6]. If, in addition, $\operatorname{span}\Delta_{K}=2$, then $g(K)=1$, hence Wang’s theorem [46] implies $K$ cannot have purely cosmetic surgeries.
@@
Test 2.5 (Jones derivative test).
@@
Test 2.6 (TQFT test).
Given a knot $K\subset S^{3}$, with diagram $D=D(K)$, compute the diagram Turaev genus $g_{T}(D)$. Next, compute the Jones polynomial $V_{K}(t)$ and its evaluation $V_{K}(e^{2\pi i/5})$. If $g_{T}(D)\leq 15$ and $V_{K}(e^{2\pi i/5})\neq 1$, then Detcherry’s theorem [10] implies that $K$ cannot have purely cosmetic surgeries.
We remark that Detcherry’s theorem [10] relies on Theorem 2.1. In addition, note that $g_{T}(D)\leq\frac{1}{2}c(D)$ for the crossing number $c(D)$, so $g_{T}(D)\leq 15$ for all diagrams up to $31$ crossings. This makes Test 2.6 quite powerful in practice.
The final triple of tests uses invariants derived from the knot Floer homology $\widehat{HFK}(K)$.
Test 2.7 (Genus 1 test).
Given a knot $K\subset S^{3}$, compute the Seifert genus $g(K)$ using $\widehat{HFK}(K)$. If $g(K)=1$, then Wang’s theorem [46] implies that $K$ cannot have purely cosmetic surgeries.
In the special case where $K$ is alternating, we have $g(K)=\frac{1}{2}\operatorname{span}\Delta_{K}$, so this test can be applied without computing $\widehat{HFK}(K)$.
Test 2.8 ($\tau$ invariant test).
@@
Test 2.9 (Hanselman HFK test).
Given a knot $K\subset S^{3}$, compute $th(K)$ and $g(K)$ using $\widehat{HFK}(K)$. If $th(K)<2g(K)(g(K)-2)$, then Hanselman’s Theorem 2.1 implies $K$ cannot have purely cosmetic surgeries.
In the special case where $K$ is alternating, this test combined with Test 2.7 provides exactly the same information as Test 2.4.
@@ 2.1. Effectiveness of the tests
To assess the comparative effectiveness of the above tests, we ran each of them on the 1,701,935 prime knots up to 16 crossings. Here is how they performed:
@@
Test 2.3: Casson invariant 1,513,776 188,159 11.1%
Test 2.4: Quick genus-thickness 1,676,441 25,494 1.5%
Test 2.5: Jones derivative 1,623,086 78,849 4.6%
Test 2.6: TQFT 1,701,883 52 0.003%
Test 2.7: Seifert genus 1 72 1,701,863 99.996%
Test 2.8: $\tau$ invariant 1,252,518 449,417 26.4%
Test 2.9: Hanselman HFK 1,694,443 7,492 0.44%
One immediate conclusion is that Test 2.6 (TQFT) is very effective. Indeed, its effectiveness seems to grow as the crossing number increases: there are over 8,000,000 knots with 17 crossings, and Detcherry has computed that only 45 of these knots have the property that $V_{K}(e^{2\pi i/5})=1$, which makes the test inconclusive [10]. Test 2.9 (Hanselman HFK) is also quite powerful. The combination of Tests  2.4, 2.5, and 2.6 is very quick to run, and serves to rule out the vast majority of knots.
@@
In the following table, $B_{n}$ represents the event “a random prime knot of up to 16 crossings survives after applying test $n$.” Then, the $(i,j)$ entry of the table is the normalized likelihood $\frac{P(B_{i}|B_{j})}{P(B_{i})}$. Thus values less than $1$ suggest that tests $i$ and $j$ are negatively correlated, while values greater than $1$ suggest that they are positively correlated. Diagonal entries represent the inverse probability $\frac{1}{P(B_{i})}$ that a knot survives after test $i$; as expected, the diagonal entries for Test 2.6 and Test 2.9 are particularly large because those tests are particularly effective.
None of the entries in the table is substantially below $1$, meaning that any negative correlations are very slight. On the other hand, there are some notable positive correlations. For instance, Test 2.4 (quick genus–thickness) is strongly positively correlated with Test 2.9 (Hanselman HFK); indeed, apart from the 72 knots of genus $1$, any knot that survives Test 2.9 must also survive Test 2.4. The strong positive correlation between Test 2.6 (TQFT) and Test 2.9 (Hanselman HFK) is more surprising, and merits further study.
Test 2.3 Test 2.4 Test 2.5 Test 2.6 Test 2.8 Test 2.9
Test 2.3 9.045 1.227 4.063 1.217 1.406 0.946
Test 2.4 66.75 1.396 12.83 1.815 66.41
Test 2.5 21.58 0.830 1.596 0.939
Test 2.6 32729 0.728 69.89
Test 2.8 3.786 1.697
Test 2.9 227.1
@@ 2.2. Results
Burton has compiled an enumeration of all 352,152,252 prime nontrivial knots up to 19 crossings [5]. We ran Tests 2.32.6 on this list of knots. These tests rule out purely cosmetic surgeries on all but two of the knots: the knots $K_{1}$ and $K_{2}$ that are shown in Figure 1. Both exceptional knots satisfy $th(K_{i})=0$, $g(K_{i})=2$, and $\tau(K_{i})=0$, so the $\widehat{HFK}$–based Tests 2.72.9 do not provide any information.
@@
Figure 1. The two prime knots up to 19 crossings that were not ruled out by Tests 2.32.9. Both exceptional knots have 18 crossings.
At the same time, Theorem 2.1 says that the only potential purely cosmetic pairs of surgery slopes on $K_{i}$ are {±$\{\pm 1\}$ and {±$\{\pm 2\}$. In each case, the filled manifolds are hyperbolic and are distinguished using verified hyperbolic volume. ∎
@@
There are good theoretical reasons why the above tests fail on $K_{1}$ and $K_{2}$. Since $K_{1}$ and $K_{2}$ are thin (meaning $th(K_{i})=0$) and have Alexander polynomials of the form appearing in [18, Theorem 5], Heegaard Floer invariants cannot distinguish the {±$\{\pm 1\}$ and {±$\{\pm 2\}$ Dehn surgeries on these knots. The Alexander polynomial is similarly powerless.
In addition, performing ±$\pm 1/5$ surgery on a crossing circle about the large twist region of $K_{i}$ has the effect of replacing 9 half-twists by $-1$ half-twist, or 10 half-twists by 0. This surgery transforms each $K_{i}$ into the unknot, while leaving the value $J_{K}(e^{2\pi i/5})$ unchanged. Compare [10, Lemma 4.2]. Thus Test 2.6 cannot provide any information about $K_{i}$.
Finally, the alternating knot $K_{1}$ belongs to one of the families excluded by Ichihara and Jong’s result [25, Theorem 1.1].
@@
Running Tests 2.32.6 on a database of 352 million knots took approximately 5 months of wall time.
@@ 3.1. Invariants based on homology
Given a closed 3-manifold $M$, consider a finite-index subgroup $H<\pi_{1}M$. The normal core $K=\operatorname{Core}(H)$ is the intersection of all conjugates of $H$. Consider the $4$–tuple
@@
In practice, computing the normal core of a subgroup $H$ requires working in Gap, which is slower than the cover enumeration in SnapPy. However, this computation is practical and the invariant $\mathcal{D}(M,n)$ seems to be powerful. In prior work, Dunfield successfully used this invariant (with $n=6$) to distinguish roughly 300,000 closed hyperbolic 3-manifolds [11, Proof of Theorem 1.4].
@@
Let $N$ be a compact oriented 3-manifold with torus boundary. The homological longitude of $N$ is a primitive homology class in $H_{1}(\partial N,{\mathbb{Z}})$ that is trivial in $H_{1}(N,{\mathbb{Q}})$. By the “half lives, half dies” lemma [20, Lemma 3.5], the homological longitude always exists and is unique up to sign.
@@
Let $N$ be a compact oriented 3-manifold with torus boundary. Let ${\lambda_{\rm hom}}$ be its homological longitude, and let ${\mu_{\rm hom}}$ any curve on $\partial N$ whose intersection number with ${\lambda_{\rm hom}}$ is $1$. Then there is an integer $k=k(N)$ such that the following hold for every relatively prime pair $(p,q)$, giving slope $s=p{\mu_{\rm hom}}+q{\lambda_{\rm hom}}$.
@@
Let $i\colon\partial N\to N$ be the inclusion map, and let $i_{*}\colon H_{1}(\partial M)\to H_{1}(M)$ be the induced homomorphism. By the “half lives, half dies” lemma [20, Lemma 3.5] and the definitions of ${\lambda_{\rm hom}}$ and ${\mu_{\rm hom}}$, we have that $i_{*}({\lambda_{\rm hom}})$ is finite order and $i_{*}({\mu_{\rm hom}})$ is infinite order in $H_{1}(N)$. Let $F$ be a maximal finite subgroup of $H_{1}(N)$ containing $i_{*}({\lambda_{\rm hom}})$. Let $C=\langle\nu\rangle$ be a maximal cyclic subgroup of $H_{1}(N)$ containing $i_{*}({\mu_{\rm hom}})$, so that $i_{*}({\mu_{\rm hom}})=m\nu$. By the maximality of $C$ and $F$, we have a splitting
@@
Now, let $s=p{\mu_{\rm hom}}+q{\lambda_{\rm hom}}$ be a Dehn filling slope. To compute the homology groups of $N(s)$, we first attach a $2$–handle $D^{2}$ along the slope $s$, and then attach the remaining 3-handle. The 3-handle has no effect on $H_{1}(N(s))$. The relevant portion of the Mayer–Vietoris sequence in reduced homology is
[…]
@@
[…]
@@ 3.2. Linearized Hodgson–Kerckhoff bounds
The next two subsections are fairly technical. The main mathematical upshot is to establish certain estimates that constrain the set of slopes that need to be analyzed for excluding cosmetic surgeries. One of the key results, Theorem 3.14, constrains the set of Dehn fillings that satisfy an upper bound on volume. A second key result, Theorem 3.19, provides an improvement on [15, Theorem 7.29]. These results are combined in Theorem 3.24, which constrains the set of pairs of slopes to check to a finite and explicitly computable set.
The results below are definitely useful; even small improvements on previous estimates provide valuable algorithmic speedups that matter a great deal when dealing with hundreds of thousands of examples. However, both the results and their proofs are technical in nature. Accordingly, the casual reader who is mainly interested in the procedural and algorithmic aspects of this work is warmly invited to jump ahead to Section 4.
@@
Note that ${\operatorname{haze}}^{-1}(h)$ is defined and monotonically decreasing on $[0,1.0196]$. See [15, Remark 4.23].
@@
As above, let $K=3.3957$ and $z_{0}=\sqrt{-2+\sqrt{5}}$. On the interval $[z_{0},1]$, define
$u(z)=K\,\frac{z^{2}(z^{4}+4z^{2}-1)}{2\,(z^{2}+1)^{3}}\quad\>\>\text{and}\quad\>\>v(z)=\frac{K}{4}\left(\frac{-2z^{5}+z^{4}-z^{3}+2z^{2}-z+1}{(z^{2}+1)^{2}}+\arctan(z)-\frac{\pi}{4}\right)\!.$
[…]
@@
[…]
@@
The conclusion that $M=N({\mathbf{s}})$ is hyperbolic and $\Sigma$ consists of geodesics is [22, Theorem 1.1]. The estimates on $\operatorname{vol}(M)$ and $\operatorname{len}(\Sigma)$ are a partial restatement of [22, Theorem 5.12], with simplified notation. Hodgson and Kerckhoff formulated their theorem in terms of functions
@@
that are defined on [22, page 1079]. The integrand in their upper bound on $\operatorname{vol}(N)-\operatorname{vol}(M)$ simplifies to our function $u(z)$. We remark that [22, Theorem 5.12] also includes lower bounds on $\operatorname{vol}(N)-\operatorname{vol}(M)$ and $\operatorname{len}(\Sigma)$, which we will not need. ∎
@@
The subscript in ${f_{\textup{pl}}}^{-1}$ stands for “piecewise linear.”
@@
The function $f(z)$ is decreasing on its domain by [2, Lemma 5.4], and concave down on $[0.83,1]$ by [2, Lemma 5.9]. Consequently, for any sub-interval $[a,b]\subset[0.83,1]$, the graph of $f$ lies above the secant line segment connecting the points $(a,f(a))$ and $(b,f(b))$. Since $f$ is decreasing, the graph of $f$ also lies to the right of this secant line.
[…]
@@
[…]
@@
[…]
Here, the first inequality uses Theorem 3.10. The second inequality uses the lower bound on ${\hat{z}}$ in previous displayed equation, combined with the fact that ${\operatorname{haze}}$ is decreasing. This proves half of the upper bound in (3.15).
@@
[…]
Here, the first inequality is the previous displayed equation. The second inequality is verified using interval arithmetic in Sage; see the ancillary files [13] for full details. Meanwhile, when $L=L({\mathbf{s}})\geq 11.7$, we have $L^{2}\geq 136.89$. Under this hypothesis, the estimate $\operatorname{len}(\Sigma)\leq\frac{2\pi}{L({\mathbf{s}})^{2}-14.41}$ is a consequence of [15, Lemma 6.10]; see the two displayed equations in the proof of [15, Lemma 8.4]. ∎

### v3

@@ 2.2. Practical tests
$\tfrac{1}{2}\operatorname{span}\Delta_{K}\leq g(K).$
If $g_{T}(D)<\operatorname{span}\Delta_{K}(\tfrac{1}{2}\operatorname{span}\Delta_{K}-2)$, then Lowrance’s theorem [40] implies th(K)<2g(K)(g(K)2)th(K)<2g(K)(g(K)-2), hence Hanselman’s Theorem 2.1 implies $K$ cannot have purely cosmetic surgeries.
If $g_{T}(D)=0$, then $K$ is alternating [11, Corollary 4.6]. If, in addition, $\operatorname{span}\Delta_{K}=2$, then $g(K)=1$, hence Wang’s theorem [51] implies $K$ cannot have purely cosmetic surgeries.
We remark that for knots in $S^{3}$, Test 2.5 and Test 2.6 are both strictly weaker than Test 2.4. However, since Test 2.4 was not available during our initial work on this project, both Test 2.5 and Test 2.6 are used in our proof of Theorem 2.11.
@@
Test 2.7 (Jones derivative test).
@@
Test 2.8 (TQFT test).
Given a knot $K\subset S^{3}$, with diagram $D=D(K)$, compute the diagram Turaev genus $g_{T}(D)$. Next, compute the Jones polynomial $V_{K}(t)$ and its evaluation $V_{K}(e^{2\pi i/5})$. If $g_{T}(D)\leq 15$ and $V_{K}(e^{2\pi i/5})\neq 1$, then Detcherry’s theorem [12, Theorem 1.4], combined with Theorem 2.1, implies that $K$ cannot have purely cosmetic surgeries.  [^1] If $V_{K}(e^{2\pi i/5})\neq 1$, Detcherry’s theorem [12, Theorem 1.4] implies that any cosmetic surgery pair for $K$ is of the form $\{\pm\frac{1}{5k}\}$. Thus Theorem 2.2 implies that $K$ cannot have purely cosmetic surgeries, even without any information about $g_{T}(D)$. This makes Test 2.8 simpler to state and apply, but does not affect its power for small-crossing knots because all diagrams up to 31 crossings satisfy $g_{T}(D)\leq 15$.
Detcherry’s argument [12] works by comparing Turaev–Viro invariants $TV_{r}(M)$, where $M$ is a Dehn filling of $K$ and $r\geq 5$ is a prime number. When $r=5$, the criterion can be expressed in terms of $V_{K}(e^{2\pi i/5})$, leading to a test that is both easy to state and easy to implement. We have not attempted to implement Detcherry’s more general result [12, Theorem 1.5] for larger values of $r$, in part because Test 2.8 is already stunningly effective.
The final pair of tests uses invariants derived from the knot Floer homology $\widehat{HFK}(K)$.
Test 2.9 ($\tau$ invariant test).
@@
Test 2.10 (HFK genus-thickness test).
Given a knot $K\subset S^{3}$, compute the Seifert genus $g(K)$ and the thickness $th(K)$ using $\widehat{HFK}(K)$. If $g(K)=1$, then Wang’s theorem [51] implies that $K$ cannot have purely cosmetic surgeries. If th(K)<2g(K)(g(K)2)th(K)<2g(K)(g(K)-2), then Hanselman’s Theorem 2.1 implies $K$ cannot have purely cosmetic surgeries.  [^2] If $g(K)\neq 2$, then Theorem 2.1 implies any cosmetic pair for $K$ is of the form $\{\pm\frac{1}{q}\}$, hence Theorem 2.2 implies $K$ has no cosmetic surgeries at all. Thus, in light of Theorem 2.2, it suffices to verify that $g(K)\neq 2$. This makes Test 2.10 simpler to state and to apply, but does not affect its power for small-crossing knots. This is because all knots up to 17 crossings (and probably further) satisfy $th(K)\leq 5$, hence $g(K)\neq 2$ implies $K$ has no cosmetic surgeries by [23, Corollary 3].
We remark that when $K$ is alternating, Test 2.10 provides exactly the same information as Test 2.6: both tests rule out cosmetic surgeries whenever $g(K)\neq 2$.
@@ 2.3. Effectiveness of the tests
To assess the comparative effectiveness of the above tests, we ran each of them on the 9,755,328 prime knots up to 17 crossings. Running a single test on this sample took approximately 1.5 days of wall time for the Alexander–based tests, 4 days of wall time for the Jones–based tests, and 6 days of wall time for the Heegaard Floer–based tests. Here, the term wall time refers to the clock time between starting the process (on one processor core) and the completion of the process.
Here is how the tests performed:
@@
Test 2.4: Alexander polynomial 9,752,895 2,433 0.025%
Test 2.5: Casson invariant 8,745,275 1,010,053 10.3%
Test 2.6: Quick genus-thickness 9,668,010 87,318 0.89%
Test 2.7: Jones derivative 9,313,098 442,230 4.5%
Test 2.8: TQFT 9,755,231 97 0.001%
Test 2.9: $\tau$ invariant 7,260,357 2,494,971 25.6%
Test 2.10: HFK genus-thickness 9,740,390 14,938 0.15%
One immediate conclusion is that Test 2.8 (TQFT) is extremely effective. Indeed, its effectiveness seems to grow as the crossing number increases: the proportion of prime knots with exactly $n$ crossings that satisfy $V_{K}(e^{2\pi i/5})=1$ is
$0.0035\%\text{ for }n=15,\quad 0.0017\%\text{ for }n=16,\quad 0.0006\%\text{ for }n=17.$
Since Test 2.8 is considerably slower than the Alexander–based tests, we note that the combination of Tests 2.5, 2.6, 2.7, and 2.8 is very fast to run when performed in that order, and serves to rule out the vast majority of knots. We also note that Test 2.4 (Alexander polynomial), which only became available after our initial work on this project, is both very fast and very powerful all on its own.
@@
In the following table, $B_{i}$ represents the event “a random prime knot of up to 17 crossings survives after applying test ii.” Then, the $(i,j)$ entry of the table is the normalized likelihood $\frac{P(B_{i}|B_{j})}{P(B_{i})}$. Thus values less than $1$ suggest that tests $i$ and $j$ are negatively correlated, while values greater than $1$ suggest that they are positively correlated. Diagonal entries represent the inverse probability $\frac{1}{P(B_{i})}$ that a knot survives after test $i$; as expected, the diagonal entries for Test 2.4, Test 2.8, and Test 2.10 are particularly large because those tests are particularly effective.
Test 2.4 Test 2.5 Test 2.6 Test 2.7 Test 2.8 Test 2.9 Test 2.10
Test 2.4 (Alexander) 4010 9.658 111.7 6.909 0 3.409 66.567
Test 2.5 (Casson) 9.658 1.334 3.851 1.394 1.413 0.952
Test 2.6 (Quick) 111.7 1.523 21.88 1.872 111.7
Test 2.7 (Jones deriv) 22.05 0.455 1.623 0.997
Test 2.8 (TQFT) 100570 0.846 127.9
Test 2.9 ($\tau$ invariant) 3.910 1.787
Test 2.10 (HFK genus) 653.0
The table contains a notable 0, meaning that exactly zero prime knots up to 17 crossings survive both Test 2.4 (Alexander) and Test 2.8 (TQFT). However, this is not actually evidence of negative correlation: if Test 2.4 and Test 2.8 were truly independent, we would expect $0.02$ knots to survive both, but of course the number of knots is an integer. Since none of the other entries in the table is substantially below $1$, we see that any negative correlations between the tests are very slight.
Some of the positive correlations in the table have a known cause. For instance, Test 2.4 (Alexander) is strongly positively correlated with Test 2.6 (quick genus–thickness) because any knot with $\Delta_{K}=1$ must also survive all other tests based on the Alexander polynomial. Similarly, Test 2.6 (quick genus–thickness) is strongly positively correlated with Test 2.10 (HFK genus-thickness) because any knot that survives Test 2.10 must also survive Test 2.6. The strong positive correlation between Test 2.8 (TQFT) and Test 2.10 (HFK genus-thickness) is more surprising, and merits further study.
@@ 2.4. Results
Burton has compiled an enumeration of all 352,152,252 prime nontrivial knots up to 19 crossings [6]. We ran Tests 2.52.8 on this list of knots. These tests rule out purely cosmetic surgeries on all but two of the knots: the knots $K_{1}$ and $K_{2}$ that are shown in Figure 1.
@@
[…]
Figure 1. The two prime knots up to 19 crossings that were not ruled out by Tests 2.52.10. Both exceptional knots have 18 crossings.
Now, Theorem 2.1 says that the only potential purely cosmetic pairs of surgery slopes on $K_{i}$ are $\{\pm 1\}$ and $\{\pm 2\}$. In each case, the filled manifolds are hyperbolic and are distinguished using verified hyperbolic volume. ∎
@@
There are good theoretical reasons why Tests 2.52.10 all fail on the knots $K_{1}$ and $K_{2}$. Both of these knots satisfy $g(K_{i})=2$ and $\tau(K_{i})=0$, so the $\widehat{HFK}$–based Tests 2.92.10 do not provide any information. Since the two knots are thin (meaning $th(K_{i})=0$) and have Alexander polynomials of the form appearing in [23, Theorem 5], Heegaard Floer invariants cannot distinguish the $\{\pm 1\}$ and $\{\pm 2\}$ Dehn surgeries on these knots. Consequently, Test 2.6 (a quicker and weaker version of Test 2.10) is similarly powerless.
In addition, performing $\pm 1/5$ surgery on a crossing circle about the large twist region of $K_{i}$ has the effect of replacing 9 half-twists by 1-1 half-twist, or 10 half-twists by 0. This surgery transforms each $K_{i}$ into the unknot, while leaving the value $J_{K}(e^{2\pi i/5})$ unchanged. Compare [12, Lemma 4.2]. Thus Test 2.8 cannot provide any information about $K_{i}$.
At the same time, the knots $K_{1}$ and $K_{2}$ have nontrivial Alexander polynomials, so Test 2.4 excludes purely cosmetic surgeries on this pair. We did not use Test 2.4 in the proof of Theorem 2.11, because Theorem 2.2 only became available after the first draft of this paper.
@@
Running Tests 2.52.8 on a database of 352 million knots took approximately 5 months of wall time (running the tests on two cores in parallel).
@@ 3.1. Invariants based on homology
Given a closed 3-manifold $M$, consider a finite-index subgroup $H<\pi_{1}M$. The normal core $K=\operatorname{Core}(H)$ is the intersection of all conjugates of $H$. Consider the 44–tuple
@@
In practice, computing the normal core of a subgroup $H$ requires working in GAP [22], which is slower than the cover enumeration in SnapPy. However, this computation is practical and the invariant $\mathcal{D}(M,n)$ seems to be powerful. In prior work, Dunfield successfully used this invariant (with $n=6$) to distinguish roughly 300,000 closed hyperbolic 3-manifolds [14, Proof of Theorem 1.4].
@@
Let $N$ be a compact oriented 3-manifold with torus boundary. The homological longitude of $N$ is a primitive homology class in $H_{1}(\partial N,{\mathbb{Z}})$ that is trivial in $H_{1}(N,{\mathbb{Q}})$. By the “half lives, half dies” lemma [25, Lemma 3.5], the homological longitude always exists and is unique up to sign.
@@
Let $N$ be a compact oriented 3-manifold with torus boundary. Let $\lambda_{\rm hom}$} be its homological longitude, and let $\mu_{\rm hom}$} be any curve on $\partial N$ whose intersection number with $\lambda_{\rm hom}$} is $1$. Then there is an integer $k=k(N)$ such that the following hold for every relatively prime pair $(p,q)$, giving slope $s=p{\mu_{\rm hom}}+q{\lambda_{\rm hom}}$.
@@
Let $i\colon\partial N\to N$ be the inclusion map, and let $i_{*}\colon H_{1}(\partial M)\to H_{1}(M)$ be the induced homomorphism. By the “half lives, half dies” lemma [25, Lemma 3.5] and the definitions of $\lambda_{\rm hom}$} and $\mu_{\rm hom}$}, we have that $i_{*}({\lambda_{\rm hom}})$ is finite order and $i_{*}({\mu_{\rm hom}})$ is infinite order in $H_{1}(N)$. Let $F$ be a maximal finite subgroup of $H_{1}(N)$ containing $i_{*}({\lambda_{\rm hom}})$. Let $C=\langle\nu\rangle$ be a maximal cyclic subgroup of $H_{1}(N)$ containing $i_{*}({\mu_{\rm hom}})$, so that $i_{*}({\mu_{\rm hom}})=m\nu$. By the maximality of $C$ and $F$, we have a splitting
@@
Now, let $s=p{\mu_{\rm hom}}+q{\lambda_{\rm hom}}$ be a Dehn filling slope. To compute the homology groups of $N(s)$, we first attach a 22–handle $D^{2}$ along the slope $s$, and then attach the remaining 3-handle. The 3-handle has no effect on $H_{1}(N(s))$. The relevant portion of the Mayer–Vietoris sequence in reduced homology is
$\ldots\to H_{1}(S^{1})\xrightarrow{\>\alpha\>}H_{1}(N)\oplus H_{1}(D^{2})\xrightarrow{\>\beta\>}H_{1}(N\cup D^{2})\xrightarrow{\>\partial\>}\widetilde{H}_{0}(S^{1})\to\ldots$
@@
$H_{1}(N(s))$ $\cong{\raisebox{1.99997pt}{$H_{1}(N)$}\!\!\left/\raisebox{-1.99997pt}{\!$\ker(\beta)$}\right.}$
$\cong{\raisebox{1.99997pt}{$B\oplus C\oplus F$}\!\!\left/\raisebox{-1.99997pt}{\!$\langle i_{*}(p{\mu_{\rm hom}}+q{\lambda_{\rm hom}})\rangle$}\right.}$
@@ 3.2. Linearized Hodgson–Kerckhoff bounds
The next two subsections are fairly technical. The main mathematical upshot is to establish certain estimates that constrain the set of slopes that need to be analyzed for excluding cosmetic surgeries. One of the key results, Theorem 3.14, constrains the set of Dehn fillings that satisfy an upper bound on volume. A second key result, Theorem 3.19, provides an improvement on [18, Theorem 7.29]. These results are combined in Theorem 3.24, which constrains the set of pairs of slopes to check to a finite and explicitly computable set.
The results below have an intrinsic intellectual value, because they are close to sharp. In addition, the applications of these results in Section 3.3 are quite useful: even small improvements on previous estimates provide valuable algorithmic speedups that matter a great deal when dealing with hundreds of thousands of examples. However, both the results and their proofs are technical in nature. Accordingly, the casual reader who is mainly interested in the procedural and algorithmic aspects of this work is warmly invited to jump ahead to Section 4.
@@
Note that ${\operatorname{haze}}^{-1}(h)$ is defined and monotonically decreasing on $[0,1.0196]$. See [18, Remark 4.23].
@@
As above, let $K=3.3957$ and $z_{0}=\sqrt{\sqrt{5}-2}$. On the interval $z_{0},1]$, define
$u(z)=K\,\frac{z^{2}(z^{4}+4z^{2}-1)}{2\,(z^{2}+1)^{3}}\quad\>\>\text{and}\quad\>\>v(z)=\frac{K}{4}\left(\frac{-2z^{5}+z^{4}-z^{3}+2z^{2}-z+1}{(z^{2}+1)^{2}}+\tan^{-1}(z)-\frac{\pi}{4}\right)\!.$
It is straightforward to check that $\int_{z}^{1}u(x)\,dx=v(z)$, hence $v^{\prime}(z)=-u(z)$. Furthermore, $u(z)>0$ on $(z_{0},1)$, hence $v(z)$ is strictly decreasing and satisfies $\lim_{z\to 1}v(z)=0$. Next, define
$f(z)=K(1-z)\exp\left(\int_{z}^{1}\frac{-(x^{4}+6x^{2}+4x+1)}{(x+1)(x^{2}+1)^{2}}\,dx\right)\!.$
@@
$\operatorname{vol}(N)-\operatorname{vol}(M)\leq v({\hat{z}})=\int_{{\hat{z}}}^{1}u(z)\,dz\qquad\text{and}\qquad\operatorname{len}(\Sigma)\leq\frac{{\operatorname{haze}}({\hat{z}})}{2\pi}.$
@@
The conclusion that $M=N({\mathbf{s}})$ is hyperbolic and $\Sigma$ consists of geodesics is [27, Theorem 1.1]. The estimates on $\operatorname{vol}(M)$ and $\operatorname{len}(\Sigma)$ are a partial restatement of [27, Theorem 5.12], with simplified notation. Hodgson and Kerckhoff formulated their theorem in terms of functions
@@
that are defined on [27, page 1079]. The integrand in their upper bound on $\operatorname{vol}(N)-\operatorname{vol}(M)$ simplifies to our function $u(z)$. We remark that [27, Theorem 5.12] also includes lower bounds on $\operatorname{vol}(N)-\operatorname{vol}(M)$ and $\operatorname{len}(\Sigma)$, which we will not need. ∎
@@
See Figure 2 for a graph. The subscript in ${f_{\textup{pl}}}^{-1}$ stands for “piecewise linear.”
[…]
Figure 2. The three functions appearing in (3.13).
@@
The function $f(z)$ is decreasing on its domain by [2, Lemma 5.4], and concave down on $[0.83,1]$ by [2, Lemma 5.9]. Consequently, for any sub-interval $a,b]\subset[0.83,1]$, the graph of $f$ lies above the secant line segment connecting the points $(a,f(a))$ and $(b,f(b))$. Since $f$ is decreasing, the graph of $f$ also lies to the right of this secant line.
Direct computation shows that the slope of the secant line from $(0.85,f(0.85))$ to $(1,0)$ is lower (that is, more negative) than 1/0.374-1/0.374, hence the secant line between those points lies above the function $f_{\text{lin}}(z)$, and the graph of $f(z)$ lies higher still. Similarly, direct computation shows that the slope of the secant line from $(0.85,f(0.85))$ to $(0.9,f(0.9))$ is lower (that is, more negative) than 1/0.452-1/0.452, hence $f(z)$ lies above ${f_{\textup{pl}}}(z)$ on the sub-interval $[0.85,0.9]$.
@@
$\min\left\{1.031-\frac{17.85}{L({\mathbf{s}})^{2}},\>0.9\right\}\><\>\min\left\{1.031-0.452\left(\frac{2\pi}{L({\mathbf{s}})}\right)^{\!2}\!,\>0.9\right\}\>\leq\>f^{-1}\left(\Big(\frac{2\pi}{L({\mathbf{s}})}\Big)^{\!2}\right)\>=\>{\hat{z}}({\mathbf{s}}).$
@@
$\operatorname{len}(\Sigma)\>\leq\>\frac{{\operatorname{haze}}({\hat{z}})}{2\pi}\>\leq\>\frac{{\operatorname{haze}}\Big(\!\min\big\{1.031-\tfrac{17.85}{L({\mathbf{s}})^{2}},\>0.9\big\}\Big)}{2\pi}.$
Here, the first inequality uses Theorem 3.10. The second inequality uses the lower bound on $^{\hat{z}}$ in previous displayed equation, combined with the fact that $\operatorname{haze}$} is decreasing. This proves half of the upper bound in (3.15).
@@
(3.16) $\operatorname{len}(\Sigma)\>\leq\>\frac{{\operatorname{haze}}\Big(\!\min\big\{1.031-\tfrac{17.85}{L({\mathbf{s}})^{2}},\>0.9\big\}\Big)}{2\pi}\><\>\frac{2\pi}{L({\mathbf{s}})^{2}-14.41}.$
Here, the first inequality is the previous displayed equation. The second inequality is verified using interval arithmetic in Sage; see the ancillary files [16] for full details. Meanwhile, when $L=L({\mathbf{s}})\geq 11.7$, we have $L^{2}\geq 136.89$. Under this hypothesis, the estimate $\operatorname{len}(\Sigma)\leq\frac{2\pi}{L({\mathbf{s}})^{2}-14.41}$ is a consequence of [18, Lemma 6.10]; see the two displayed equations in the proof of [18, Lemma 8.4]. ∎